The length of a rectangle is twice its width. If the perimeter of the rectangle is 36 yd, what is the area.

Answer by checkley75(3666)

L=2W
2W+2(2W)=36
2W+4W=36
6W=36
W=36/6
W=6 ANSWER FOR THE WIDTH.
L=2*6=12 ANSWER FOR THE LENGTH.
PROOF:
2*6+2*12=36
12=24=36
36=36
AREA=6*12=72 YDS^2.